OpenStudy (anonymous):

find the lim as x approaches zero of 2-2cos(3x)/3x

OpenStudy (raffle_snaffle):

do you know your limit laws?

OpenStudy (anonymous):

some what

OpenStudy (raffle_snaffle):

I guess you could apply L'Hospitals Rule. Did you learn his rule yet?

OpenStudy (aravindg):

2-2cos(3x)=2(1-cos 3x) Now apply 1-cos t formula

OpenStudy (anonymous):

no I have not. I've only learned the basics. like when the numerator exponent is higher the lim does not exist and denominator degree is higher the lim is zero

OpenStudy (aravindg):

Try what I said above.

OpenStudy (raffle_snaffle):

Listen to aravinda

OpenStudy (aravindg):

*Aravind

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I did that what's next?

OpenStudy (aravindg):

What did you get?

OpenStudy (anonymous):

2(1-COS3X)/3X from this point do I just plug in the zero?

OpenStudy (anonymous):

That'll give me 0/0 which its not allowed

OpenStudy (anonymous):

please help, I'm very clueless and I can't afford to fell that test tomorrow

OpenStudy (psymon):

Well, in general: \[\lim_{x \rightarrow 0}\frac{ 1-cosx }{ x }=0\]As long as the x angle of cosine and the x int he denominator match, you can say its 0 and NOT undefined. So it looks like you just have 2*0 = 0

OpenStudy (psymon):

@have_sabr

OpenStudy (anonymous):

That makes sense Thank you very much! :)

OpenStudy (psymon):

Yep, np :3

hero (hero):

Given: \[\lim_{x \rightarrow 0}\frac{2\cos(3x)}{3x}\] Taking the RIGHT side limit we have: \[\lim_{x \rightarrow 0^{+}}\frac{2\cos(3x)}{3x} = \infty\] Taking the LEFT side limit we have: \[\lim_{x \rightarrow 0^{-}}\frac{2\cos(3x)}{3x}= -\infty\] Therefore, the two sided limit does not exist.